OpenStudy (anonymous):

If y varies directly with x and y = -7 when x = -1, find x when y = -84. A.84 B.7 C.-12

OpenStudy (anonymous):

x=ky put y=-7 and x=-1 k=1/7 now put y=-84 x=-84/7 x=12 C

OpenStudy (anonymous):

oops x=-12 missed - sign

OpenStudy (anonymous):

Thank you very much for your help @ArslanQureshi ! :)

OpenStudy (anonymous):

np :)